### v4

@@ 2.2 Technological History and Roadmap
Table 1: Overview of commonly used datasets. Orange-marked ones are selected to evaluate different methods in Sec. 4.
@@
Deepfake Generation LFW huang2008labeled Image 10K Facial images captured under various lighting conditions, poses, expressions, and occlusions at 250$\times$250 resolution.
@@
CelebA liu2015deep Image 200K The dataset includes over 200K facial images from more than 10K individuals, each with 40 attribute labels.
@@
VGGFace parkhi2015deep Image 2600K A super-large-scale facial dataset involving a staggering 26K participants, encompassing a total of 2600K facial images.
@@
VoxCeleb1 nagrani2017voxceleb Video 100K voices A large scale audio-visual dataset of human speech, the audio includes noisy background interference.
@@
CelebA-HQ karras2018progressive Image 30K A high-resolution facial dataset consisting of 30K face images, each with a resolution of 1024$\times$1024 resolution.
@@
VGGFace2 cao2018vggface2 Image 3000K A large-scale facial dataset, expanded with greater diversity in terms of ethnicity and pose compared to VGGFace.
@@
VoxCeleb2 chung2018voxceleb2 Video 2K hours A dataset that is five times larger in scale than VoxCeleb1 and improves racial diversity.
@@
FFHQ karras2019style Image 70K The dataset with over 70K high-resolution (1024$\times$1024) facial images, showcasing diverse ethnicity, age, and backgrounds.
@@
MEAD wang2020mead Video 40 hours An emotional audiovisual dataset provides facial expression information during conversations with various emotional labels.
@@
CelebV-HQ zhu2022celebv Video 35K videos The dataset’s video clips have a resolution of more than 512$\times$512 resolution and are annotated with rich attribute labels.
@@
Forgery Detection DeepfakeTIMIT korshunov2018deepfakes Audio-Video 640 videos The dataset is evenly divided into two versions: LQ (64$\times$64) and HQ (128$\times$128), with all videos using face swapping forgery.
@@
FF++ rossler2019faceforensics++ Video 6K videos Comprising 1K original videos and manipulated videos generated using five different forgery methods.
@@
DFDCP dolhansky2019deepfake Audio-Video 5K videos The preliminary dataset for The Deepfake Detection Challenge includes two face-swapping methods.
@@
DFD DfD Video 3K videos The dataset comprises 3K deepfake videos generated using five forgery methods.
@@
Deeperforensics jiang2020deeperforensics Video 60K videos The videos feature faces with diverse skin tones, and rich environmental diversity was considered during the filming process.
@@
DFDC dolhansky2020deepfake Audio-Video 128K videos The official dataset for The Deepfake Detection Challenge and contain a substantial amount of interference information.
@@
Celeb-DF li2020celeb Video 1K videos Comprising 408 genuine videos from diverse age groups, ethnicities, and genders, along with 795 DeepFake videos.
@@
Celeb-DFv2 li2020celeb Video 6K videos An expanded version of Celeb-DFv1, this dataset not only increases in quantity but also the diversity.
@@
FakeAVCeleb khalid2021fakeavceleb Audio-Video 20K videos A novel audio-visual multimodal deepfake detection dataset, deepfake videos generated using four forgery methods.
$\bullet$ Discriminative Neural Network. Convolutional Neural Networks (CNNs) lecun1998gradient ; krizhevsky2012imagenet ; he2016deep ; liu2022convnet ; emo ; eatformer have played a pivotal role in the history of deep learning. LeNet lecun1998gradient , as the pioneer of CNNs, showcased the charm of machine learning. AlexNet krizhevsky2012imagenet and ResNet he2016deep made deep learning feasible. Recently, ConvNeXt liu2022convnet has achieved excellent results surpassing those of Swin-Transformer liu2021swin . The Transformer architecture initially proposed vaswani2017attention in 2017. The core idea involves using self-attention mechanisms to capture dependencies between different positions in the input sequence, enabling global modeling of sequences. ViT dosovitskiy2020image demonstrates that using Transformer in the field of computer vision can still achieve excellent performance and PVT wang2021pyramid overcomes the challenges of adapting Transformer to various dense prediction tasks. In addition, Swin-Transformer liu2021swin addresses the limitations of Transformer in handling high-resolution image processing tasks. Subsequently, Swin-Transformer V2 liu2022swin further improves the model’s efficiency and the resolution of manageable inputs.
$\bullet$ Neural Radiance Field (NeRF) is first introduced in 2020 mildenhall2020nerf , with its core idea revolving around the use of volume rendering and implicit neural fields to represent and reconstruct both geometric and illumination information of 3D scenes gao2022nerf . Compared to traditional 3D methods, it exhibits higher visual quality and is currently widely applied in tasks such as 3D geometry enhancement sharma2023neural ; irshad2023neo , segmentation liu2023instance and 6D pose estimation min2023neurocs . In addition, Some notable works hwang2023faceclipnerf ; yang2023urbangiraffe combining NeRF as a supplement to 3D information and generation models are particularly prominent at present.
2.3 Datasets, Metrics, and Losses
$\bullet$ Dataset. Given the various datasets in surveyed fields, we use numerical labels to save post-textual space. 1) Commonly used deepfake generation datasets include LFW huang2008labeled , CelebA liu2015deep , CelebA-HQ karras2018progressive , VGGFace parkhi2015deep , VGGFace2 cao2018vggface2 , FFHQ karras2019style , Multi-PIE moore2010multi , VoxCeleb1 nagrani2017voxceleb , VoxCeleb2 chung2018voxceleb2 , MEAD wang2020mead , MM CelebA-HQ xia2021tedigan , CelebAText-HQ sun2021multi , CelebV-HQ zhu2022celebv , TalkingHead-1KH wang2021one , LRS2 afouras2018deep , LRS3 afouras2018lrs3 , etc. 2) Commonly used forgery detection datasets include UADFV li2018ictu , DeepfakeTIMIT korshunov2018deepfakes , FF++ rossler2019faceforensics++ , Deeperforensics-1.0 jiang2020deeperforensics , DFDCP dolhansky2019deepfake , DFDC dolhansky2020deepfake , Celeb-DF li2020celeb , Celeb-DFv2 li2020celeb , FakeAVCeleb khalid2021fakeavceleb , DFD DfD , WildDeepfake zi2020wilddeepfake , KoDF kwon2021kodf , UADFV yang2019exposing , Deephy narayan2022deephy , DF-Platter narayan2023df , etc. We summarize popular datasets in Table 1 .
$\bullet$ Metric. 1) For deepfake generation tasks, commonly used metrics include: Peak Signal-to-Noise Ratio (PSNR) wang2023memory , Structured Similarity (SSIM) wang2004image , Learned Perceptual Image Patch Similarity (LPIPS) zhang2018unreasonable , Fréchet Inception Distance (FID) heusel2017gans , Kernel Inception Distance (KID) binkowski2018demystifying , Cosine Similarity (CSIM) wang2021one , Identity Retrieval Rate (ID Ret) wang2018cosface , Expression Error chaudhuri2019joint , Pose Error ruiz2018fine , Landmark Distance (LMD) around the mouths chen2018lip , Lip-sync Confidence (LSE-C) prajwal2020lip , Lip-sync Distance (LSE-D) prajwal2020lip , etc. 2) For forgery generation commonly uses: Area Under the ROC Curve (AUC) le2023quality , Accuracy (ACC) wang2022m2tr , Equal Error Rate (EER) huang2023implicit , Average Precision (AP) tan2023learning , F1-Score chen2021image , etc. Detailed definitions are explained in Sec. 4.1.
$\bullet$ Loss Function. VAE-based approaches generally employ reconstruction loss and KL divergence loss kingma2014auto . Commonly used reconstruction loss functions include Mean Squared Error, Cross-Entropy, LPIPS zhang2018unreasonable , and perceptual johnson2016perceptual losses. GAN-based methods further introduce adversarial loss goodfellow2014generative to increase image authenticity, while diffusion-based works introduce denoising loss function ho2020denoising .
2.4 Related Research Domains
@@
$\bullet$ Head Swapping replaces the entire head information rather than only the face of the target image shu2022few , including facial contours and hairstyle, with that of the source image. However, in terms of facial attributes, Head Swap only replaces identity attributes, while leaving other attribute information unchanged. Recently, some methods wang2023hs ; han2023generalist based on Diffusion have been proposed and have shown promising results in achieving good performance.
@@
$\bullet$ Face Super-resolution aims to enhance the resolution of low-resolution face images to generate high-resolution face images scsnet ; moser2023hitchhiker . This task is closely related to various deepfake sub-tasks. In the early stages of Face Swapping natsume2018rsgan ; sun2018hybrid and Talking Face Generation fan2015photo ; chen2019hierarchical , there is consistently a low-resolution issue in the generated images and videos. This problem is addressed by incorporating face super-resolution methods into the models mir2023dit ; xu2022high to improve output quality. In terms of technical approaches, FSR can be categorized into methods based on CNNs huang2017wavelet ; chen2019sequential ; chen2020learning , GANs KoMulti ; moser2024diffusion , reinforcement learning shi2019face , and ensemble learning jiang2019atmfn .
@@
$\bullet$ Face Reconstruction refers to the process of recreating the three-dimensional appearance of an individual’s face based on one or multiple 2D facial images morales2021survey ; sharma20223d . Facial reconstruction often plays an intermediate role in various deepfake sub-tasks. In Face Swapping and Face Reenactment tasks, facial parameters are reconstructed using 3DMM, and the model’s parameters can be controlled directionally. Additionally, reconstructing a 3D facial model is one of the methods to address the issue of facial artifacts in videos generated under large facial poses. Common technical approaches for facial reconstruction include methods based on 3DMM maninchedda2017fast ; lin2020towards , epipolar geometry feng20183d , one-shot learning tulsiani2017multi ; xing2019self , shadow shape reconstruction kemelmacher20103d ; jiang20183d , and hybrid learning-based reconstruction dou2017end ; chaudhuri2020personalized .
@@
$\bullet$ Face Inpainting, a.k.a. face completion, aims to reconstruct missing regions in face images caused by external factors such as occlusion and lighting while preserving facial texture information is crucial in this process zhang2023image . This task is a crucial sub-task of image inpainting, and the current methods are mostly based on deep learning that can be roughly divided into two categories: GAN based yildirim2023diverse ; zhou2023superior and Diffusion based xu2024personalized ; yang2024pgdiff .
@@
$\bullet$ Body Animation aims to alter the entire bodily pose while unchanging the overall body information yu2023bidirectionally . The goal is to achieve a modification of the target image’s entire body posture using an optional driving image or video, aligning the body of the target image with the information from the driving signal. The mainstream implementation path for body animation is based on GANs zhou2022cross ; zhao2022thin , and Diffusion wang2023leo ; ma2023follow ; wang2023disco ; xu2023magicanimate .
@@
$\bullet$ Portrait Style Transfer aims to reconstruct the style of a target image to match that of a source image by learning the stylistic features of the source image yang2022vtoonify ; peng2023portraitbooth . The goal is to preserve the content information of the target image while adapting its style to that of the source image cai2023image . Common applications include image cross-domain style transfer, such as transforming real face images into animated face styles perez2024styleavatar ; feng20243d . Methods based on GANs abdal20233davatargan ; xu2022your3demoji ; jiang2023scenimefy and Diffusion liu2023portrait ; hur2024expanding have achieved high-quality performance in this task.
@@
$\bullet$ Makeup Transfer aims to achieve style transfer learning from a source image to a target image jiang2020psgan ; li2023hybrid . Existing models have achieved initial success in applying and removing makeup on target images chang2018pairedcyclegan ; gu2019ladn ; liu2021psgan++ ; li2023hybrid , allowing for quantitative control over the intensity of makeup. However, they perform poorly in transferring extreme styles zhong2023sara ; yang2022elegant ; yan2023beautyrec . Existing mainstream methods are based on GANs sun2023ssat ; hao2022cumtgan ; jiang2020psgan .
@@
$\bullet$ Adversarial Sample Detection focuses on identifying whether the input data is an adversarial sample han2023interpreting . If recognized as such, the model can refuse to provide services for it, such as throwing an error or not producing an output yuan2021current . Current deepfake detection models often rely on a single cue from the generation process as the basis for detection, making them vulnerable to specific adversarial samples. Furthermore, relatively little work has focused on adversarial sample testing in terms of model generalization capability and detection evaluation.
3 Deepfake Inspections: A Survey
This section systematically examines four generation tasks: Face Swapping (Sec. 3.1.1), Talking Face Generation (Sec. 3.1.3), Face Reenactment (Sec. 3.1.2), Facial Attribute Editing (Sec. 3.1.4). Additionally, we review techniques for forgery detection (commonly known as deepfake detection) in Sec. 3.2. We document and compare the most representative works in the main text. Furthermore, we discuss several related domains that garnered significant attention in Sec. 3.3.
3.1 Deepfake Generation
@@
Table 2: Overview of representative face swapping methods. Notations: ➊ Self-build, ➋ CelebA-HQ, ➌ FFHQ, ➍ VGGFace2, ➎ VGGFace, ➏ CelebV, ➐ CelebA, ➑ VoxCeleb2, ➒ LFW, ➓ KoDF. Abbreviations: SIGGRAPH (SIG.), EUROGRAPHICS (EG.), GANs (G.), VAEs (V.), Diffusion (D.), Split-up and Integration (SI.).
@@
Traditional Graphics Blanz et alblanz2004exchanging EG.’04 3DMM Manual intervention, unnatural output. Early face-swapping efforts simplified manual interaction steps.
@@
Bitouk et albitouk2008face SIG.’08 SI. Manual intervention, attribute loss. A three-phase implementation framework with the help of a pre-constructed face database to match faces that are similar to the source face in terms of posture and lighting.
@@
Sunkavalli et alsunkavalli2010multi TOG’10 SI. Poor generalizability, frequent artifacts. Early work on face exchange was realized using image processing methods such as smooth histogram matching technique.
@@
Dale et aldale2011video SIG.’11 3DMM Poor generalizability and output quality. Early work on face exchange, proposing an improved Poisson mixing approach to achieve face swapping in video through frame-by-frame face replacement.
@@
Lin et allin2012face ICME’12 zhu2009unsupervised 3DMM Poor generalizability, frequent artifacts. An attempt to construct a personalized 3D head model to solve the artifact problem occurring in face swapping in large poses.
@@
Mosaddegh et almosaddegh2015photorealistic ACCV’14 gross2010multi milborrow2010muct SI. Poor generalizability and output quality. A diverse form of face swapping where facial components can be targeted for replacement.
@@
Nirkin et alnirkin2018face FG’18 burgos2013robust SI. Poor generalization ability and resolution. Transfer of expressions and poses by building some 3D variable models and training facial segmentation networks to maintain target facial occlusion.
@@
Generative Adversarial Network IPGAN bao2018towards CVPR’18 guo2016ms G.+V. Poor output image quality, frequent artifacts. Using two encoders to encode facial identity and attribute information separately for facial information decoupling and swapping.
@@
RSGAN natsume2018rsgan SIG.’18 G.+V. Loss of lighting information. Using two independent VAE modules to represent the latent spaces of the face and hair regions, respecti -vely, with the replacement of identity information in the latent space implemented.
@@
Sun et alsun2018hybrid ECCV’18 zhang2015beyond G.+3DMM Poor ability to preserve face feature attributes. Implementing in two stages: the first stage involves replacing the identity information of the face reg -ion, while the second stage achieves complete facial rendering.
@@
FSGAN nirkin2019fsgan ICCV’19 maze2018iarpa G. Poor ability to preserve face feature attributes. Two novel loss functions are introduced to refine the stitching in the face fusion phase following the swapping process.
@@
FaceShifter li2019faceshifter CVPR’20 ➊➋➌➎ G. Poor ability to preserve face feature attributes. Face swapping is realized in two stages, the firstly AEI-Net improves the output image quality level, and the second HEAR-Net is targeted to focus on abnormal regions for image recovery.
@@
Zhu et alzhu2020deepfakes AAAI’20 G.+V. Inability to process facial contour information. First show of the applicability of deepfake to keypoint invariant de-identification work.
DeepFaceLab liu2023deepfacelab PR’20 G.+V. Poor output image resolution. A mature and complete framework for face swapping.
@@
SimSwap chen2020simswap MM’20 ➊➍ G.+V. Poor ability to preserve face feature attributes. ID modules and weak feature matching loss functions are proposed to find a balance between identity information replacement and attribute information retention.
@@
MegaFS zhu2021one CVPR’21 ➊➋➌ G. Poor ability to preserve face feature attributes. The first method allows for face swapping on images with a resolution of one million pixels.
FaceInpainter li2021faceinpainter CVPR’21 ➊➋➌➎ G.+3DMM Poor representation of image details. A two-stage framework innovatively implements heterogeneous domains face swapping.
@@
HifiFace wang2021hififace IJCAI’21 G.+3DMM Uses a large number of parameters. A 3D shape-aware identity extractor is proposed to achieve better retention of attribute information such as facial shape.
@@
FSGANv2 nirkin2022fsganv2 TPAMI’22 maze2018iarpa G. Unable to process posture differences effectively. An extension of the FSGAN method that combines Poisson optimization with perceptual loss enhances the output image facial details.
StyleSwap xu2022styleswap ECCV’22 ➊➎➑ G. Unable to process posture differences effectively. Proposing an exchange-guided ID reversal strategy to enhance the performance of attribute informa -tion replacement during face exchange.
RAFSwap xu2022region CVPR’22 G. Unable to process posture differences effectively. Local facial region awareness branch with global source feature adaptation (SFA) branch is proposed to better achieve the preservation of target image attribute information.
@@
FSLSD xu2022high CVPR’22 ➊➋ G. Poor ability to preserve face feature attributes. Potential semantic de-entanglement is realized to obtain facial structural attributes and appearance attributes in a hierarchical manner.
@@
Kim et alkim2022smooth CVPR’22 ➌➍ G. Unable to process posture differences effectively. An identity embedder is proposed to enhance the training speed under supervision.
@@
3DSwap li20233d CVPR’23 ➋➌ G.+3DMM Unable to process posture differences effectively. A 3d-aware approach to the face-swapping task, de-entangling identity and attribute features in latent space to achieve identity replacement and attribute feature retention.
FALCO barattin2023attribute CVPR’23 ➋➒ G. Poor ability to handle facial occlusion. Oriented with privacy-preserving applications, the method directly employs latent space of pre-trained GANs to achieve the identity of anonymized images while preserving facial attributes.
WSC-Swap ren2023reinforced ICCV’23 ➊➋➌➎ G.+3DMM Poor resolution of the output image . Two mutually independent encoders are proposed to encode attribute information outside the face region and semantic-level non-identical facial attributes inside the face region.

### v5

@@ 2.3. Datasets, Metrics, and Losses
Table 1. Overview of commonly used datasets. Orange-marked are selected to evaluate different methods.
@@
Deepfake Generation LFW (huang2008labeled, ) Image 10K Facial images captured under various lighting conditions, poses, expressions, and occlusions at $\times250$ resolution.
@@
CelebA (liu2015deep, ) Image 200K The dataset includes over 200K facial images from more than 10K individuals, each with 40 attribute labels.
@@
VGGFace (parkhi2015deep, ) Image 2600K A super-large-scale facial dataset involving a staggering 26K participants, encompassing a total of 2600K facial images.
@@
VoxCeleb1 (nagrani2017voxceleb, ) Video 100K voices A large scale audio-visual dataset of human speech, the audio includes noisy background interference.
@@
CelebA-HQ (karras2018progressive, ) Image 30K A high-resolution facial dataset consisting of 30K face images, each with a resolution of $\times1024$ resolution.
@@
VGGFace2 (cao2018vggface2, ) Image 3000K A large-scale facial dataset, expanded with greater diversity in terms of ethnicity and pose compared to VGGFace.
@@
VoxCeleb2 (chung2018voxceleb2, ) Video 2K hours A dataset that is five times larger in scale than VoxCeleb1 and improves racial diversity.
@@
FFHQ (karras2019style, ) Image 70K The dataset with over 70K high-resolution ($\times1024$) facial images, showcasing diverse ethnicity, age, and backgrounds.
@@
MEAD (wang2020mead, ) Video 40 hours An emotional audiovisual dataset provides facial expression information during conversations with various emotional labels.
@@
CelebV-HQ (zhu2022celebv, ) Video 35K videos The dataset’s video clips have a resolution of more than $\times512$ resolution and are annotated with rich attribute labels.
@@
Forgery Detection DeepfakeTIMIT (korshunov2018deepfakes, ) Audio-Video 640 videos The dataset is evenly divided into two versions: LQ ($\times64$) and HQ ($\times128$), with all videos using face swapping forgery.
@@
FF++ (rossler2019faceforensics++, ) Video 6K videos Comprising 1K original videos and manipulated videos generated using five different forgery methods.
@@
DFDCP (dolhansky2019deepfake, ) Audio-Video 5K videos The preliminary dataset for The Deepfake Detection Challenge includes two face-swapping methods.
@@
DFD (DfD, ) Video 3K videos The dataset comprises 3K deepfake videos generated using five forgery methods.
@@
Deeperforensics (jiang2020deeperforensics, ) Video 60K videos The videos feature faces with diverse skin tones, and rich environmental diversity was considered during the filming process.
@@
DFDC (dolhansky2020deepfake, ) Audio-Video 128K videos The official dataset for The Deepfake Detection Challenge and contain a substantial amount of interference information.
@@
Celeb-DF (li2020celeb, ) Video 1K videos Comprising 408 genuine videos from diverse age groups, ethnicities, and genders, along with 795 DeepFake videos.
@@
Celeb-DFv2 (li2020celeb, ) Video 6K videos An expanded version of Celeb-DFv1, this dataset not only increases in quantity but also the diversity.
@@
FakeAVCeleb (khalid2021fakeavceleb, ) Audio-Video 20K videos A novel audio-visual multimodal deepfake detection dataset, deepfake videos generated using four forgery methods.
2.4. Related Research Domains
@@
$\bullet$ Head Swapping. This task replaces the entire head region of a target image, including facial contours and hairstyle, with that of a source image (shu2022few, ). Despite the larger replacement area, only identity-related attributes are transferred, while other facial attributes remain unchanged. Recently, diffusion-based methods (han2023generalist, ) have emerged and demonstrated promising performance.
@@
$\bullet$ Face Super-resolution. This task enhances low-resolution images to produce high-resolution outputs (scsnet, ; moser2023hitchhiker, ). It is closely related to many deepfake sub-tasks. Early methods in Face Swapping (natsume2018rsgan, ; sun2018hybrid, ) and Talking Face Generation (fan2015photo, ; chen2019hierarchical, ) often suffered from low-resolution outputs in synthesized images and videos. This limitation has been alleviated by integrating face super-resolution modules into generative pipelines (mir2023dit, ; xu2022high, ), significantly improving visual quality. Technically, face super-resolution approaches can be categorized into CNNs (huang2017wavelet, ; chen2019sequential, ; chen2020learning, ), GANs (KoMulti, ; moser2024diffusion, ), reinforcement learning (shi2019face, ), and ensemble learning (jiang2019atmfn, ).
@@
$\bullet$ Face Reconstruction. This task refers to reconstructing the three-dimensional facial structure of an individual from one or multiple 2D images (morales2021survey, ; sharma20223d, ). Facial reconstruction often serves as an intermediate step in various deepfake sub-tasks. In Face Swapping and Face Reenactment, 3DMM are widely used to recover facial parameters, enabling controllable identity or expression manipulation. Reconstructing a 3D facial model also helps mitigate artifacts that occur in synthesized videos under large pose variations. Technical approaches to face reconstruction include 3DMM (maninchedda2017fast, ; lin2020towards, ), epipolar-geometry (feng20183d, ), one-shot learning (tulsiani2017multi, ; xing2019self, ), shadow shape reconstruction (kemelmacher20103d, ; jiang20183d, ), and hybrid learning‑based reconstruction (dou2017end, ; chaudhuri2020personalized, ).
@@
$\bullet$ Face Inpainting. This task aims to reconstruct missing regions in face images caused by external factors such as occlusion and lighting while preserving facial texture information is crucial in this process (zhang2023image, ). This task is a crucial sub-task of image inpainting, and the current methods are mostly based on deep learning that can be roughly divided into two categories: GAN based (yildirim2023diverse, ; zhou2023superior, ) and Diffusion based (xu2024personalized, ; yang2024pgdiff, ).
@@
$\bullet$ Body Animation. This task aims to alter the entire bodily pose while unchanging the overall body information (yu2023bidirectionally, ). The goal is to achieve a modification of the target image’s entire body posture using an optional driving image or video, aligning the body of the target image with the information from the driving signal. The mainstream implementation path for body animation is based on GANs (zhou2022cross, ; zhao2022thin, ), and Diffusion (wang2023leo, ; ma2023follow, ; wang2023disco, ; xu2023magicanimate, ).
@@
$\bullet$ Portrait Style Transfer. This task aims to reconstruct the style of a target image to match that of a source image by learning the stylistic features of the source image (yang2022vtoonify, ; peng2023portraitbooth, ). The goal is to preserve the content information of the target image while adapting its style to that of the source image (cai2023image, ). Common applications include image cross-domain style transfer, such as transforming real face images into animated face styles (perez2024styleavatar, ; feng20243d, ). Methods based on GANs (abdal20233davatargan, ; xu2022your3demoji, ; jiang2023scenimefy, ) and Diffusion (liu2023portrait, ; hur2024expanding, ) have achieved high-quality performance in this task.
@@
$\bullet$ Makeup Transfer. This task aims to achieve style transfer learning from a source image to a target image (jiang2020psgan, ; li2023hybrid, ). Existing models have achieved initial success in applying and removing makeup on target images (chang2018pairedcyclegan, ; gu2019ladn, ; liu2021psgan++, ; li2023hybrid, ), allowing for quantitative control over the intensity of makeup. However, they perform poorly in transferring extreme styles (zhong2023sara, ; yang2022elegant, ; yan2023beautyrec, ). Existing mainstream methods are based on GANs (hao2022cumtgan, ; jiang2020psgan, ).
@@
$\bullet$ Adversarial Sample Detection. This task focuses on identifying whether the input data is an adversarial sample (han2023interpreting, ). If recognized as such, the model can refuse to provide services for it, such as throwing an error or not producing an output (yuan2021current, ). Current deepfake detection models often rely on a single cue from the generation process as the basis for detection, making them vulnerable to specific adversarial samples. Furthermore, relatively little work has focused on adversarial sample testing in terms of model generalization capability and detection evaluation.
2.5. Applications of Deepfake Detection
Deepfake detection is essential for ensuring the integrity and security of digital media as synthetic content becomes increasingly realistic. Several major institutions have deployed detection tools in practical settings. Youtube incorporates automated deepfake classifiers into its content moderation pipeline to identify manipulated media before large-scale dissemination (youtube, ). Financial institutions such as the Bank of China (Chinese_BANK, ) employ deepfake-oriented liveness detection to prevent video-based impersonation fraud. These real-world deployments highlight the essential role of deepfake detection in protecting individuals, financial systems, and public information integrity.
2.6. Ethical and Societal Considerations
The rapid development of deepfake technologies has raised major ethical and societal concerns, particularly regarding privacy invasion, identity misuse, and the spread of manipulated content. In response, several countries and regions have introduced regulatory frameworks for synthetic media. The European Union incorporates transparency requirements into the AI Act (Europe, ) and the Digital Services Act (Europe2, ), mandating clear labeling of manipulated content and stricter platform obligations. China’s “Provisions on the Administration of Deep Synthesis Internet Information Services” establish comprehensive rules on watermarking, traceability, identity verification, and platform accountability (China, ). These policies illustrate a global movement toward responsible governance of synthetic media, emphasizing transparency and protection against misuse.
3. Deepfake Inspections: A Survey
3.1. Deepfake Generation
@@
Table 2. Overview of representative face swapping methods. Notations: ➊ Self-build, ➋ CelebA-HQ, ➌ FFHQ, ➍ VGGFace2, ➎ VGGFace, ➏ CelebV, ➐ CelebA, ➑ VoxCeleb2, ➒ LFW, ➓ KoDF. Abbreviations: SIGGRAPH (SIG.), EUROGRAPHICS (EG.), GANs (G.), VAEs (V.), Diffusion (D.), Split-up and Integration (SI.).
@@
Traditional Graphics Blanz et al(blanz2004exchanging, ) EG.’04 3DMM Manual intervention, unnatural output. Early face-swapping efforts simplified manual interaction steps.
@@
Bitouk et al(bitouk2008face, ) SIG.’08 SI. Manual intervention, attribute loss. A three-phase implementation framework with the help of a pre-constructed face database to match faces that are similar to the source face in terms of posture and lighting.
@@
Sunkavalli et al(sunkavalli2010multi, ) TOG’10 SI. Poor generalizability, frequent artifacts. Early work on face exchange was realized using image processing methods such as smooth histogram matching technique.
@@
Dale et al(dale2011video, ) SIG.’11 3DMM Poor generalizability and output quality. Early work on face exchange, proposing an improved Poisson mixing approach to achieve face swapping in video through frame-by-frame face replacement.
@@
Lin et al(lin2012face, ) ICME’12 (zhu2009unsupervised, ) 3DMM Poor generalizability, frequent artifacts. An attempt to construct a personalized 3D head model to solve the artifact problem occurring in face swapping in large poses.
@@
Mosaddegh et al(mosaddegh2015photorealistic, ) ACCV’14 (gross2010multi, )(milborrow2010muct, ) SI. Poor generalizability and output quality. A diverse form of face swapping where facial components can be targeted for replacement.
@@
Nirkin et al(nirkin2018face, ) FG’18 (burgos2013robust, ) SI. Poor generalization ability and resolution. Transfer of expressions and poses by building some 3D variable models and training facial segmentation networks to maintain target facial occlusion.
@@
Generative Adversarial Network IPGAN (bao2018towards, ) CVPR’18 (guo2016ms, ) G.+V. Poor output image quality, frequent artifacts. Using two encoders to encode facial identity and attribute information separately for facial information decoupling and swapping.
@@
RSGAN (natsume2018rsgan, ) SIG.’18 G.+V. Loss of lighting information. Using two independent VAE modules to represent the latent spaces of the face and hair regions, respecti -vely, with the replacement of identity information in the latent space implemented.
@@
Sun et al(sun2018hybrid, ) ECCV’18 (zhang2015beyond, ) G.+3DMM Poor ability to preserve face feature attributes. Implementing in two stages: the first stage involves replacing the identity information of the face reg -ion, while the second stage achieves complete facial rendering.
@@
FSGAN (nirkin2019fsgan, ) ICCV’19 (maze2018iarpa, ) G. Poor ability to preserve face feature attributes. Two novel loss functions are introduced to refine the stitching in the face fusion phase following the swapping process.
@@
FaceShifter (li2019faceshifter, ) CVPR’20 ➊➋➌➎ G. Poor ability to preserve face feature attributes. Face swapping is realized in two stages, the firstly AEI-Net improves the output image quality level, and the second HEAR-Net is targeted to focus on abnormal regions for image recovery.
@@
Zhu et al(zhu2020deepfakes, ) AAAI’20 G.+V. Inability to process facial contour information. First show of the applicability of deepfake to keypoint invariant de-identification work.
@@
SimSwap (chen2020simswap, ) MM’20 ➊➍ G.+V. Poor ability to preserve face feature attributes. ID modules and weak feature matching loss functions are proposed to find a balance between identity information replacement and attribute information retention.
@@
MegaFS (zhu2021one, ) CVPR’21 ➊➋➌ G. Poor ability to preserve face feature attributes. The first method allows for face swapping on images with a resolution of one million pixels.
@@
HifiFace (wang2021hififace, ) IJCAI’21 G.+3DMM Uses a large number of parameters. A 3D shape-aware identity extractor is proposed to achieve better retention of attribute information such as facial shape.
@@
FSGANv2 (nirkin2022fsganv2, ) TPAMI’22 (maze2018iarpa, ) G. Unable to process posture differences effectively. An extension of the FSGAN method that combines Poisson optimization with perceptual loss enhances the output image facial details.
@@
FSLSD (xu2022high, ) CVPR’22 ➊➋ G. Poor ability to preserve face feature attributes. Potential semantic de-entanglement is realized to obtain facial structural attributes and appearance attributes in a hierarchical manner.
@@
Kim et al(kim2022smooth, ) CVPR’22 ➌➍ G. Unable to process posture differences effectively. An identity embedder is proposed to enhance the training speed under supervision.
@@
3DSwap (li20233d, ) CVPR’23 ➋➌ G.+3DMM Unable to process posture differences effectively. A 3d-aware approach to the face-swapping task, de-entangling identity and attribute features in latent space to achieve identity replacement and attribute feature retention.